OpenStudy (anonymous):

Please help!-Celine has a bottle that contains 20% of milk and the rest water. The bottle has one liter of water.Write an equation in one variable that can be used to find the total number of liters of milk and water in the bottle.

OpenStudy (anonymous):

you already know how many liters of water are in the bottle. you are told it is 1 now you need an expression to find how many liters of milk are in the bottle

OpenStudy (anonymous):

since 1 is 80% of the total, the total is \(\frac{1}{.8}\) and since 20% of that is milk, the amount of milk is \[.2\times \frac{1}{.8}\]

OpenStudy (anonymous):

which is not hard to compute, since \[.2\times \frac{1}{8}=\frac{.2}{.8}=\frac{2}{8}=\frac{1}{4}\]

OpenStudy (anonymous):

Ok, Thank you :)